# Deployment and validation of predictive 6-dimensional beam diagnostics through generative reconstruction with standard accelerator elements

### v2

@@ I.5 Reconstructed 6-dimensional beam phase space
Finally, Fig. 4 illustrates the reconstructed 6-dimensional phase space of the beam at the entrance to the Q1 magnet. Figure. 4a shows 2-dimensional projections of the beam distribution, providing a comprehensive view of the correlations among all beam coordinate axes, while Fig. 4b shows a subset of 3-dimensional projections along coordinate axes. The reconstruction predicts non-linear tails in the transverse dimensions and accurately reproduces the nonlinear energy chirp, as observed by direct measurements shown in Fig. 3
@@
$\epsilon_{nx}$ 0.42 0.40 (0.42$\pm$0.01) mm mrad
$\epsilon_{ny}$ 0.39 0.40 (0.39$\pm$0.01) mm mrad
$\beta_{x}$ 4.53 4.67 (4.59$\pm$0.15) m
$\beta_{y}$ 7.78 8.12 (7.87$\pm$0.29) m
$\alpha_{x}$ $-0.56$ $-0.50$ ($-0.52\pm$0.04) rad
$\alpha_{y}$ 1.25 1.17 (1.16$\pm$0.09) rad
@@
$\epsilon_{nz}$ 137.34$\pm$17.24 140.05 (141.29$\pm$6.27) mm mrad
$\sigma_{z}$ $0.45\pm 0.01$ 0.49 (0.47$\pm$0.01) mm
$\sigma_{\delta}$ $0.24+0.01$ 0.22 (0.22$\pm$0.00) %
We also calculated ensemble-averaged quantities of the beam as summarized in Table 1. The 6-dimensional reconstruction accurately reproduces transverse emittance values and Twiss parameters. These predicted values have quantitative agreement within 10% of those obtained by 4-dimensional phase space reconstructions using the GPSR analysis procedure outlined in [40]. For the longitudinal phase space at the end of the beamline, the reconstructed values agrees well with conventional calculations of longitudinal bunch properties, such as the RMS longitudinal bunch length, momentum deviation, and emittance (where uncertainty estimates come from systematic uncertainties in image thresholding, see details in Methods). Finally, the RMS slice energy spread--a critical quantity of interest for XFELs--is predicted to be 0.01%, which is consistent with typical values observed during normal operations at the PAL-XFEL.
In addition, we investigated the uncertainty by performing multiple reconstruction runs under different configurations of the number of iterations and data samples for the training. This analysis aims to evaluate the robustness and reproducibility of the methodology with respect to random initialization of neural networks and update of the particle distribution via gradient-based optimizations. We considered 3 additional cases: i) training with all 144 datasets for 2,400 iterations, ii) training with the same datasets for 1,200 iterations, and iii) training with 80 datasets using different XLIN phase settings for 1,000 iterations. The reconstructed beam parameters listed in Table 1 in parentheses indicates the averaged values with standard deviation $1\sigma$. It shows that the reconstruction methodology provides well converged values, exhibiting low sensitivity to random seeding variations. Furthermore, we found that the predicted beam distributions for those cases also remain in qualitative agreement with the experimental measurements.
@@ II Discussions
We have validated prediction accuracy of the GPSR methodology using standard accelerator elements, including the accelerating cavity and spectrometer magnet, without relying on specialized deflecting structures to extract $(z,\delta)$ correlations. The reconstructed 6-dimensional phase space not only shows excellent agreement with the training and test datasets, but also remains in quantitatively and qualitatively agreement with measured beam distributions further downstream in the beamline, which are excluded from the reconstruction training. These observations provides strong evidence that the GPSR can be effectively applied to systems that establish transverse–longitudinal correlations, regardless of the specific source, thereby greatly enhancing its flexibility and accessibility for beam diagnostics. For example, in low-energy proton and heavy-ion beam facilities, buncher cavities have been utilized with specialized bunch shape monitor (BSMs) for longitudinal emittance measurements [57, 58, 59]. It is thus expected that the reconstruction methodology, combined with systematic phase scan of the buncher cavity at the spectrometer screen and quadrupole scans, can provide not only the longitudinal phase space, but also the complete 6-dimensional phase space of the beam, without requiring the use of BSMs.
@@
The reconstruction methodology offers broad applicability across accelerator science. In particular, the information of high-fidelity beam phase space can be used with non-differentiable simulations that consider collective and nonlinear effects. This compatibility extends its utility to a wide range of challenges in accelerator development, including beamline optimization, predictive beam characterization along multiple beamline location with forward-backward particle trackings, and model calibration for future digital twin frameworks [60, 61]. In the context of FEL facilities, the reconstructed phase space can be applied to the measurement-based studies on suppression of emittance growth induced by the CSR effect and investigation of matching conditions along the beamline, thus to eventually enhance the FEL quality. Moreover, in advanced accelerator concepts where beam quality preservation is one of the most important tasks [62, 63], the optimization of conditions for injection into wakefields (either based on structures or plasmas) becomes critically important. The present framework provides a powerful diagnostic tool to support such optimization, thereby enhancing the performance of the future advanced acceleration schemes.
@@
In case of the uniqueness of the beam phase space through the reconstruction process, we will perform additional dedicated investigations on uncertainty quantification to further enhance the accuracy and reliability of the reconstruction methodology.
@@
The GPSR methodology presented in this work paves the way for significant advances in the realization of uncertainty-aware beam diagnostics with broader applicability in a variety of accelerator facilities with different particle species, including accelerator-based light sources, proton and heavy-ion accelerators, advanced accelerator concepts, and various test facilities.
@@ III.1 PAL-XFEL
The PAL-XFEL facility consists of a photoinjector, 3 bunch compressor sections, and number of accelerator columns. The photoinjector comprises 2.856 GHz, 1.6 cell copper-cathode RF gun and two 3.0 m accelerating cavities [64, 65]. Wavelength of the UV laser illuminating the cathode is 257 nm. UV laser distribution is basically Gaussian shape in both transverse and longitudinal planes, but truncated in the transverse plane to obtain low emittance against the space charge force along the injector section. A full width at half maximum UV length was 2.8 ps. Nominal beam charge is set to 250 pC during normal operation for user service. When the beam is generated along the photoinjector, it is propagating through the laser heater and Hard X-ray Linac 1 (HL1) section as shown in Fig. 1 of Ref. [10]. The main purpose of the laser heater is to increase the slice energy spread of the beam and suppress a micro-bunching instability. The HL1 section consists of four 3.0 m accelerating cavities, with two of them interconnected as single module. Typical kinetic energy of the beam after the HL1 is approximately 380 MeV, but we turned off the second accelerating module in this demonstration to only use the XLIN as the accelerating cavity. The kinetic energy of the beam becomes reduced to 260.5 MeV.
@@ III.2 Setup for data acquisition and GPSR training
In the demonstration section, operational phase and length of the XLIN cavity is 11.424 GHz and 0.6 m. The bending angle and arc length of the bunch compressor dipole magnet are $4.97$ deg and 0.2 m, respectively. The distance between Q1 magnet and YAG1 screen is about 12.6 m. In case of the downstream validation section, the operational frequency of 1.066 m S-band TCAV is 2.856 GHz, where the spectrometer dipole magnet has the bending angle and arc length of 15 deg and 0.7 m.
We set the range of the Q1 magnet strength from $-2.77$ to $3.16$ T/m with 16 samples. We also set the XLIN cavity phase from $-228.5$ to $-193.5$ deg with 5-degree steps including XLIN turned off. The XLIN voltage used in the differentiable simulations is 8.4 MV/m, estimated by the beam energy measurement at the YAG2 screen. During the data acquisition (DAQ), the Q2 magnet was fixed at a strength of $+2.54$ T/m while Q34343-43 - 4 were turned off. This setting ensured that the beam size at the center of the Q1 scan range was minimized in both horizontal and vertical planes, thus forming quadratic response of the beam envelope. In addition, under these conditions, we measured the fractional energy spread distribution while minimizing the horizontal beam size contribution [66].
In the GPSR process, the initial phase space distribution was sampled from a multivariate normal distribution where the scaling factor was set to $10^{-3}$ to ensure that the initial 6-dimensional phase space is on the order of mm in position and mrad in momentum. We used fully connected neural networks with two hidden layers connected by Tanh activation functions, where each layer contains 20 neurons. The back-propagation, based on the gradient-descent method where the learning rate was set to 0.01, was performed using a mean absolute error-based loss function estimated by the comparison between experimental measurements and predictions [44].
@@ III.3 Beam measurements in dispersive section
During measurements in dispersive section for the reconstruction model training, a current, corresponding to the magnetic field of the BC magnets, was fixed. The change of the beam energy induced by the XLIN cavity with respect to the RF phase results in not only the evolution of the beam shape but also the shift of the beam centroid position at the YAG1 screen. Accordingly, we set the procedure for the experimental data acquisition as follows: i) we first found the phase of the XLIN that provides the minimum horizontal beam size at the YAG1 screen. This is the case where the energy chirp is minimized. ii) We then set the BC dipole current to adjust the beam position to be at the center of the YAG1 screen; this is a procedure of defining the reference case when the XLIN is turned on. At the same time, iii) we fixed the BC dipole magnetic field for data acquisition. Finally, iv) we set the XLIN phase range that provides the beam images that are positioned within the given aspect ratio of the YAG1 screen.
@@ III.4 Image processing and downstream measurements
Due to the low SNR of the YAG1 image, we suspected that some of the beam distributions were unintentionally cut off during the image processing. In contrast, the SNR of the YAG2 image was typically larger than 100. As a result, the measured LPS and $(y,\delta)$ correlations without threshold subtraction led to underestimation of the predicted distributions. Therefore, we applied finite level of the threshold value to the YAG2 image to attenuate the beam signal and compare it with the prediction. When the BC was turned off, the threshold level of the LPS and $(y,\delta)$ images were set to approximately $10-12.5$% of the maximum intensity of the image. Similarly, in the case when the BC was turned on, the threshold level of the LPS image was set to roughly 22.5% of the maximum intensity. The error of the longitudinal emittance was calculated using 3 cases with different threshold levels; $\pm$2.5based on the reference threshold value of 10%.
@@ IV Data availability
The data presented in this paper and other findings of this study are available from the corresponding author upon reasonable request.

### v3

@@ I.5 Reconstructed 6-dimensional beam phase space
Finally, Fig. 4 illustrates the reconstructed 6-dimensional phase space of the beam at the entrance to the Q1 magnet. Figure. 4a shows 2-dimensional projections of the beam distribution, providing a comprehensive view of the correlations among all beam coordinate axes, while Fig. 4b shows a subset of 3-dimensional projections along coordinate axes. The reconstruction predicts non-linear tails in the transverse dimensions and accurately reproduces the nonlinear energy chirp, as observed by direct measurements shown in Fig. 3. In case of the phase space regarding $\delta$, the center position is slightly shifted by 0.07%, which corresponds to 0.2 MeV energy deviation compared to the reference energy of 260.5 MeV (See Methods on how to determine the reference energy). Therefore, This reconstruction suggests that more accurate reconstruction could be achieved by applying a slight adjustment to the reference energy of the beam.
@@
[…]
$\alpha_{x}$ 0.56-0.56 0.50-0.50 ($52\pm0.0$4) rad
[…]
@@
[…]
$\sigma_{\delta}$ $0.24+0.01$ 0.22 ($2±\pm0$.00) %
We also calculated ensemble-averaged quantities of the beam as summarized in Table 1. These predicted values of the 6-dimensional reconstruction are quantitatively consistent within 10% of those obtained by 4-dimensional phase space reconstructions using the GPSR analysis procedure outlined in Ref.[22]. For the longitudinal phase space at the end of the beamline, the reconstructed values agrees well with conventional calculations of longitudinal bunch properties, such as the RMS longitudinal bunch length, momentum deviation, and emittance (where uncertainty estimates come from systematic uncertainties in image thresholding, see details in Methods). Finally, the RMS slice energy spread-a critical quantity of interest for XFELs-is predicted to be 0.01%, which is consistent with typical values observed during normal operations at the PAL-XFEL.
In addition, we investigated the uncertainty of the reconstructed phase space by performing multiple training runs under different configurations of the number of iterations and data samples. We considered 3 cases additionally: i) training with all 144 datasets for 2,400 iterations, ii) training with the same datasets for 1,200 iterations, and iii) training with 80 datasets using different XLIN phase settings for 1,000 iterations. The reconstructed beam parameters listed in Table 1 in parentheses indicates the averaged values with standard deviation $1\sigma$. In addition, contour lines in Fig. 4a show specific percentiles of the phase space distributions. In the third case, correlations regarding $x$ are somewhat deviated compared to the other cases, implying that the chosen cavity phases may not have been optimal to get beam correlations during the training. Nevertheless, the overall results show that the reconstruction methodology provides well converged values, exhibiting low sensitivity to random seeding variations. Furthermore, we found that the predicted beam distributions for those cases also remain in qualitative agreement with the experimental measurements.
@@ II Discussions
We have validated prediction accuracy of the GPSR methodology using standard accelerator elements, including the accelerating cavity and spectrometer magnet, without relying on specialized deflecting structures to extract $(z,\delta)$ correlations. The reconstructed 6-dimensional phase space not only shows excellent agreement with the training and test datasets, but also remains in quantitatively and qualitatively agreement with measured beam distributions further downstream in the beamline, which are excluded from the reconstruction training. These observations provides strong evidence that the GPSR can be effectively applied to systems that establish transverse–longitudinal correlations, regardless of the specific source, thereby greatly enhancing its flexibility and accessibility for beam diagnostics.
@@
The reconstruction methodology offers broad applicability across accelerator science. In particular, the information of high-fidelity beam phase space can be used with non-differentiable simulations that consider collective and nonlinear effects. This compatibility extends its utility to a wide range of challenges in accelerator development, including beamline optimization, predictive beam characterization along multiple beamline location with forward-backward particle trackings, and model calibration for future digital twin frameworks [52]. In the context of FEL facilities, the reconstructed phase space can be applied to the measurement-based studies on suppression of emittance growth induced by the CSR effect and investigation of matching conditions along the beamline, thus to eventually enhance the FEL quality. Moreover, in advanced accelerator concepts where beam quality preservation is one of the most important tasks [53, 54], the optimization of conditions for injection into wakefields (either based on structures or plasmas) becomes critically important. The present framework provides a powerful diagnostic tool to support such optimization, thereby enhancing the performance of the future advanced acceleration schemes.
@@
In order to apply the GPSR method for the low-energy beam, consideration of space charge effect is necessary. Recently, implementation of the space charge kick matrix into the differentiable simulation is under development [55]. We will further investigate reconstruction for space charge-dominated beams. In addition, in the case of the uniqueness of the beam phase space through the reconstruction process, we will perform additional dedicated investigations to enhance the accuracy and reliability of the reconstruction methodology.
@@
The GPSR methodology presented in this work paves the way for significant advances in the realization of predictive beam diagnostics using readily available elements in a variety of accelerator facilities such as accelerator-based light sources, centers for the advanced accelerator concepts, and beam test facilities.
@@ III.1 PAL-XFEL
The PAL-XFEL facility consists of a photoinjector, 3 bunch compressor sections, and number of accelerator columns. The photoinjector comprises 2.856 GHz, 1.6 cell copper-cathode RF gun and two 3.0 m accelerating cavities [56]. Wavelength of the UV laser illuminating the cathode is 257 nm. UV laser distribution is basically Gaussian shape in both transverse and longitudinal planes, but truncated in the transverse plane to obtain low emittance against the space charge force along the injector section. A full width at half maximum UV length was 2.8 ps. Nominal beam charge is set to 250 pC during normal operation for user service. When the beam is generated along the photoinjector, it is propagating through the laser heater and Hard X-ray Linac 1 (HL1) section as shown in Fig. 1 of Ref. [6]. The main purpose of the laser heater is to increase the slice energy spread of the beam and suppress a micro-bunching instability. The HL1 section consists of four 3.0 m accelerating cavities, with two of them interconnected as single module. In this demonstration, we turned off the second accelerating module to only use the XLIN as the accelerating cavity. The kinetic energy of the beam becomes reduced to 260.5 MeV.
@@ III.2 Setup for data acquisition and GPSR training
In the demonstration section, operational phase and length of the XLIN cavity is 11.424 GHz and 0.6 m. The bending angle and arc length of the bunch compressor dipole magnet are $4.97$ deg and 0.2 m, respectively. The distance between Q1 magnet and screen 1 is about 12.6 m. In case of the downstream validation section, the operational frequency of 1.066 m S-band TCAV is 2.856 GHz, where the spectrometer dipole magnet has the bending angle and arc length of 15 deg and 0.7 m.
We set the range of the Q1 magnet strength from 2.77-2.77 to $3.16$ T/m with 16 samples to get the nominal parabolic beam envelope used for the conventional quadrupole scan. We also set the XLIN cavity phase from 228.5-228.5 to 193.5-193.5 deg with 5-degree steps including XLIN turned off. In this range of the cavity phase, changes of the energy chirp were clearly captured as reported in Ref. [44]. The initial beam energy before modulation by the XLIN was measured at the screen 2 using the spectrometer magnet. Likewise, the energy change due to the XLIN cavity was measured at the same screen. While we referred to these values as the initial condition of the GPSR process, we found that the reconstruction suggests the reference beam energy and XLIN voltage to be 260.5 MeV and 8.4 MV/m, respectively. In this setup, the reconstructed phase space is obtained where the center of each phase space is placed close to zero (e.g., $\braket{\delta}\sim 0$).
In case of the dipole magnet modeling, a linear fringe field integral (FINT) was used to consider the fringe field [57]. Whereas, a hard-edge field in the quadrupole magnet was adapted since the fringe field effect is negligible in our setup. During the data acquisition (DAQ), the Q2 magnet was fixed at a strength of +2.54+2.54 T/m while Q343-4 were turned off. This setting ensured that the beam size at the center of the Q1 scan range was minimized in both horizontal and vertical planes, thus forming quadratic response of the beam envelope. In addition, under these conditions, we measured the fractional energy spread distribution while minimizing the horizontal beam size contribution [58].
In the GPSR process, the initial phase space distribution was sampled from a multivariate normal distribution where the scaling factor was set to 0.001 to ensure that the initial 6-dimensional phase space is on the order of mm in position and mrad in momentum. We used fully connected neural networks with two hidden layers connected by Tanh activation functions, where each layer contains 20 neurons. The back-propagation, based on the gradient-descent method where the learning rate was set to 0.01, was performed using a mean absolute error-based loss function estimated by the comparison between experimental measurements and predictions [39].
@@ III.3 Beam measurements in dispersive section
During measurements in dispersive section for the reconstruction model training, a current, corresponding to the magnetic field of the BC magnets, was fixed. The change of the beam energy induced by the XLIN cavity with respect to the RF phase results in not only the evolution of the beam shape but also the shift of the beam centroid position at the screen 1. Accordingly, we set the procedure for the experimental data acquisition as follows: i) we first found the phase of the XLIN that provides the minimum horizontal beam size at the screen 1. This is the case where the energy chirp is minimized. ii) We then set the BC dipole current to adjust the beam position to be at the center of the screen 1; this is a procedure of defining the reference case when the XLIN is turned on. At the same time, iii) we fixed the BC dipole magnetic field for data acquisition. Finally, iv) we set the XLIN phase range that provides the beam images that are positioned within the given aspect ratio of the screen 1.
@@ III.4 Image processing and downstream measurements
Due to the low SNR of the screen 1 image, we suspected that some of the beam distributions were unintentionally cut off during the image processing. In contrast, the SNR of the screen 2 image was typically larger than 100. As a result, the measured LPS and $(y,\delta)$ correlations without threshold subtraction led to underestimation of the predicted distributions. Therefore, we applied finite level of the threshold value to the screen 2 image to attenuate the beam signal and compare it with the prediction. When the BC was turned off, the threshold level of the LPS and $(y,\delta)$ images were set to approximately 1012.510-12.5% of the maximum intensity of the image. Similarly, in the case when the BC was turned on, the threshold level of the LPS image was set to roughly 22.5% of the maximum intensity. The error of the longitudinal emittance was calculated using 3 cases with different threshold levels; $\pm$2.5% based on the reference threshold value of 10%.
@@ IV Data availability
Software and example of simulated demonstration are available at [60]. The experimental data presented in this paper can be made available from the corresponding authors upon reasonable request.